OpenStudy (anonymous):

write a recursive definition for each sequence.-2,7,16,25

OpenStudy (anonymous):

help me please @whpalmer4

jimthompson5910 (jim_thompson5910):

pick any term and subtract the previous term what difference do you get?

OpenStudy (anonymous):

positive 9

jimthompson5910 (jim_thompson5910):

so we're adding 9 to each term to get the next term

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

-2 + 9 = 7 7+9 = 16 16+9 = 25

jimthompson5910 (jim_thompson5910):

How do you generalize this?

OpenStudy (anonymous):

im not sure. are you asking for a equation?

jimthompson5910 (jim_thompson5910):

what's the rule to get the next term?

jimthompson5910 (jim_thompson5910):

how do I find the term after 25?

OpenStudy (anonymous):

you will add 9 to 25.

jimthompson5910 (jim_thompson5910):

so say you want to find the nth term, what's the rule to do this?

OpenStudy (anonymous):

will it be n+9

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

actually it will be \[\Large a_{n-1} + 9\] where \(\Large a_{n-1}\) is the term just before the nth term

jimthompson5910 (jim_thompson5910):

so \[\Large a_{n} = a_{n-1} + 9\]

jimthompson5910 (jim_thompson5910):

This basically says nth term = (term right before the nth term) + 9

jimthompson5910 (jim_thompson5910):

and that's how recursion works in this case

OpenStudy (anonymous):

i dont know what that is??? o.0

jimthompson5910 (jim_thompson5910):

you've never seen that notation before?

jimthompson5910 (jim_thompson5910):

maybe you've seen it as a(n) and a(n-1) ??

OpenStudy (anonymous):

ok that looks more like what i know. yes i do know what it is.

jimthompson5910 (jim_thompson5910):

ok so it would be a(n) = a(n-1) + 9 which translates to nth term = (term right before the nth term) + 9

jimthompson5910 (jim_thompson5910):

So if you wanted to know the 55th term for instance, then you would have to know the 54th term. To know the 54th term, you'd have to know the 53rd term, etc etc. So you can see that it's a lot of work using recursion.

OpenStudy (anonymous):

yes i do. i get it now thank you for you time :D

jimthompson5910 (jim_thompson5910):

you're welcome